# Differentially Private Synthetic Data via Foundation Model APIs 2: Text

### v1

@@ C.9 Aug-PE Convergence under One Private Sample
Real deal Mexican food in Chadds Ford, PA. Excellent burritos, tacos, and tamales. I plan to be a regular. Meals served traditionally with sliced radishes, pumpkin seeds and lime. Freshly prepared fillings, onion, cilantro, and red and green sauces. Brightly colored building across from Bavarian Motors. Nothing fancy, but great authentic food.
iter 0
@@
I recently visited this restaurant having heard that they had one of the best taco stands in town. I was not disappointed! The tacos were incredibly delicious and tasted like they were made with fresh ingredients. The portions were generous, and the prices were extremely reasonable. The service was excellent too, with the staff being attentive and accommodating. The ambience of the restaurant was also great, making it a perfect spot for a casual night out. Overall, I had a fantastic experience and would definitely recommend this place to all those who love tacos. A definite 5-star rating!
iter 1
@@
Despite having heard that they had one of the best taco stands in town, I was not disappointed. The tacos were incredibly delicious and tasted like they were made with fresh ingredients. The portions were generous, and the prices were reasonable. The staff was attentive and accommodating. The ambience was great, making it a perfect spot for a casual night out. Overall, I had a fantastic experience and would highly recommend this place to all those craving Mexican cuisine. A solid 5-star rating from me!
iter 2
@@
Despite having heard that they had some of the best Mexican food in town, the crowded parking lot stands in contrast. However, that was not the case. The food tasted authentic and fresh, with the guacamole and salsa being particularly noteworthy. The portions were generous, and the prices were reasonable. The staff was attentive and accommodating. The ambience was great, making it a perfect spot for a date night or a family dinner. Overall, I had a fantastic experience and would highly recommend this
iter 3
@@
Many had heard that they had amazing Mexican food in town, but the crowded parking lot stands in contrast. However, once you step inside, the experience is incredible. The food is authentic, and the guacamole and salsa being particularly noteworthy. Prices were as well quite reasonable. The staff was attentive and accommodating. The ambiance was great, making it perfect for a date night or family dinner. I had a fantastic experience and highly recommend the restaurant to anyone seeking quality Mexican cuisine.
iter 4
@@
Many locals had heard that they had the best Mexican food in town and the hype stands in its truth. However, upon stepping into the experience, it was incredible. The food was fresh, flavorful, and authentic with the guacamole and salsa being particularly noteworthy. The portions were well-sized and satisfying. The staff was attentive and accommodating. The ambiance was cozy and intimate, making it perfect for a romantic night out or casual dinner with friends. I had a great time and highly
iter 5
@@
Many locals had raved about the best authentic Mexican food in town and they were not exaggerating. Walking in to the restaurant was refreshing. The food was fresh, the guacamole and salsa were noteworthy. The margaritas were strong and satisfying. The staff were attentive and accommodating. The ambiance was cozy and intimate, making it perfect for a romantic dinner with loved ones. The prices were great and the portions were generous. The fajitas were sizzling and the tacos were packed with flavor. Overall, this
iter 6
@@
Many locals have raved about the best authentic Mexican food in town and they were not exaggerating. Walking in the restaurant was refreshing. The decor was colorful, lively and festive. The margarita was perfectly balanced and satisfying. The waiter was attentive and accommodating. The ambiance was cozy and intimate, making it perfect for a romantic dinner. The chips and salsa were packed with flavor. Overall, the food was great and the portions were generous. The fajitas were sizzling and the chicken was tender. The guacam
iter 7
@@
Many locals consistently rave about the amazing authentic Mexican cuisine and we were not disappointed. The flavors were refreshing, the decor colorful and lively, and the margaritas satisfying. The waiter was attentive and accommodating. The outdoor seating was perfect for a romantic dinner and the chips and salsa were packed with flavor. Overall, the food was top-notch and the portions were generous. The chicken enchiladas and guacamole were particularly noteworthy. It was a great experience and we highly recommend this restaurant to anyone looking for a delicious meal and a
iter 8
@@
Many foodies rave about the amazing authentic Mexican cuisine and they were not wrong. The flavors were richly robust, colorful and enticing, and the margaritas were top-notch. The enchiladas and guacamole were particularly outstanding. The service was attentive and accommodating and the outdoor seating was perfect for a leisurely dinner and people-watching. From the chips and salsa to the flavorful entrees, the food was superb with generous portions. The menu offers a large selection of delectable options which makes it easy to recommend to anyone looking for a satisfying meal
iter 9
@@
Many foodies have raved about the amazing authentic Mexican cuisine. The flavors were bold and delicious. The tacos were rich and flavorful, with enticing toppings and freshly made guacamole. The service was attentive and accommodating, and the outdoor seating was perfect for a warm summer evening. From the crispy chips and flavorful salsa to the perfectly spiced dishes, the food was generous and satisfying. The menu offers a large selection of vegetarian and meat options which makes it easy to please anyone’s taste buds. Overall, the experience was excellent and highly recommended for people looking for an authentic Mexican dining experience.
@@
The authors propose a new statistic, called Expected Perturbation Score (EPS), for adversarial detection. Based on EPS, the authors develop a Maximum Mean Difference (MMD) metric to measure the difference between test samples and natural samples, and further propose an EPS-based adversarial detection method (EPS-AD). Sufficient theoretical analysis and extensive experiments demonstrate the correctness and effectiveness of the proposed method. Strength: 1. This paper is easy to read. 2. The authors propose a new statistical method, called Expected Perturbation Score (EPS), which is able to obtain enough information to identify adversarial examples with only one example after various perturbations. 3. Sufficient theoretical analysis is performed to demonstrate that EPS is able to simulate the difference between the two distributions under mild conditions. Furthermore, extensive experimental results demonstrate the superiority of the proposed EPS-AD. 4. The proposed EPS will be an effective statistic in many applications, such as out-of-distribution detection and anomaly detection. Weaknesses: This paper is easy to read and interesting but still has some minor issues, please refer to weaknesses.
iter 0
@@
Review Example: Overall, this paper provides a comprehensive discussion on privacy issues related to machine learning applications. The authors discuss the concept of differential privacy as a potential solution to protect individual privacy while still maintaining the accuracy of the model. They also provide examples of implementation and evaluation of privacy-preserving techniques using real-world datasets. Strengths: - The paper provides a thorough explanation of differential privacy and how it can be utilized in various machine learning techniques. - The use of real-world datasets to demonstrate implementation and evaluation of privacy techniques adds credibility to the work. - The paper highlights the importance of privacy in machine learning and provides readers with potential solutions to protect individual privacy. Weaknesses: - More evaluation of the privacy techniques could be provided to better understand their effectiveness. - The paper focuses primarily on differential privacy and does not explore other privacy-preserving techniques. - The practical implementation of the techniques may be complex for those without a strong understanding of machine learning and data analysis. In summary, this paper is a strong contribution to the area of social aspects of machine learning, specifically in the field of privacy. It analyzes the concept of differential privacy, its implementation, and evaluation to protect individual privacy in machine learning models. Despite some limitations, the paper contributes significantly to the field and is recommended for acceptance.
iter 1
@@
In their paper ’Privacy in Machine Learning: An Overview’, the authors provide a comprehensive discussion on privacy issues in machine learning. The authors discuss the concept of differential privacy as a potential solution for individual privacy while maintaining the utility of the data. They also provide examples and analysis of privacy-preserving techniques on real-world datasets such as the Stanford and MNIST datasets. The thorough explanation of differential privacy and how it can be utilized in various machine learning techniques adds credibility to the paper. The paper highlights the importance of privacy in machine learning and provides a potential solution to individual privacy. Weaknesses: More evaluation of the privacy techniques could be provided to better understand their effectiveness. The paper primarily focuses on differential privacy and does not explore other privacy-preserving techniques. The technical implementation of the techniques may be difficult for those without a strong understanding of machine learning and statistical analysis. Overall, this paper contributes significantly to the area of social aspects of machine learning, and it is recommended for acceptance as a good paper.
iter 2
@@
In this paper, titled ’Policy Machine Learning: An Overview of the Discussion on Privacy Issues in Machine Learning’, the authors discuss privacy as a potential concern for individual privacy while maintaining the integrity of the data. They also provide examples of privacy-preserving techniques on real-world data from Stanford and MNIST. The thorough exploration of differential privacy can be utilized effectively which adds credibility to the paper. The paper emphasizes the importance of privacy in machine learning and provides a valuable contribution to the field. Weaknesses include the evaluation of techniques to be used to assess their effectiveness. The paper focuses on privacy issues and does not explore fairness-preserving methods. With its contribution to the social aspects of machine learning and statistical analysis, the paper is recommended with a rating of 8 as a good paper.
iter 3
@@
In their research paper, Inference Attack Policy Machine Learning: An Interpretable and Almost True Framework for Predictive Analytics, the authors highlight potential concerns for individual privacy while discussing the importance of privacy in machine learning. They also provide examples of how sensitive data from ImageNet and MNIST datasets can be utilized effectively while ensuring thorough differential privacy which adds credibility to the paper. The research emphasizes the importance of interpretability in machine learning, making a valuable contribution to the field of social aspects of machine learning. We recommend including case studies of how interpretability can be used to assess their effectiveness. The paper also outlines how it does not explore fairness and ethics methods. With this contribution to the field of machine learning and statistical modeling, the authors provide a valuable framework for policy inference attack in machine learning.
iter 4
@@
In their research paper, ’Inference Attacks in Machine Learning: An Interpretability and Almost Interpretability Framework and its Application to Privacy and Analytics’, the authors highlight the need for protecting sensitive data in machine learning. They provide examples of sensitive data from ImageNet and NIST datasets, emphasizing the importance of being thorough in privacy protection to ensure credibility to their research. The paper stresses the importance of interpretability in machine learning. By making a valuable contribution to this field, it provides case studies of how interpretability can be used to assess the effectiveness of machine learning models. The paper outlines various approaches to exploring fairness, transparency, and ethics in machine learning. The results of the study contribute to the need for a comprehensive policy to prevent inference attacks in machine learning.
iter 5
@@
In our research paper, entitled ’Adversarial Attacks in Machine Learning: An Interpretability-Almost-Explainability Framework and its Application to Private Data Analysis’, the authors emphasize the need for protecting sensitive data in machine learning. They provide examples by using data from Inet and MNIST dataset, and address the importance of privacy to ensure the credibility of the results. The paper is well-written and well-structured, making a valuable contribution to the field. Additionally, it highlights the importance of interpretability to enhance the effectiveness of machine learning models. The paper also focuses on fairness, transparency, and ethics in machine learning and the study presents a comprehensive analysis in adversarial attacks. We highly recommend accepting this good paper.
iter 6
@@
In our research paper titled ’Adversarial Attacks in Machine Learning: An Interpretable and Robustness-Enhancing Framework and Empirical Data Analysis’, the authors emphasize the significance of interpretability in machine learning. They provide a comprehensive approach using Integrated Gradients and M-Taylor expansions, to address the challenges and ensure the robustness of results. The paper is well-written, making valuable contributions to the field, and emphasizes the importance of interpretability to enhance the effectiveness of machine learning. Moreover, the study presents a comprehensive approach in defending against adversarial attacks. Therefore, I recommend accepting this good paper.
@@
In this section, we only use one private example in Alg. 1 to generate one synthetic sample. We qualitatively examine if the synthetic sample from Aug-PE increasingly resembles this specific private sample over the PE iterations. This offers a clearer illustration of Aug-PE’s convergence behavior. Specifically, at each iteration, we generate $K$ variations for the current synthetic sample, use the private sample to identify and vote for its nearest synthetic sample based on their embeddings, and select the nearest synthetic sample for the next iteration. Tb. 30 and Tb. 31 show the generations results from GPT-3.5 under one Yelp private sample and one OpenReview private sample, respectively.
@@
As shown Tb. 30, after the voting, the selected synthetic sample relates to the term “taco”, a word present in the private example. By the second iteration, the synthetic sample includes the term “Mexican food”, which aligns with the central theme of the private example. By the fifth iteration, the phrase “authentic Mexican food” surfaces in the synthetic sample, resonating with phrases like “real deal Mexican food” and “great authentic food” from the private example. This demonstrates that the synthetic sample increasingly aligns with the private sample as the iterations progress.
@@
In the OpenReview example presented in Tb. 30, we note that the initial synthetic sample at iteration 0 pertains to the privacy aspects of machine learning, whereas the private sample focuses on adversarial detection and robustness. As the iterations progress, by iteration 4, the topic of synthetic sample shifts to “inference attack in machine learning”, which aligns with the robustness theme of the private sample. By the fifth iteration, terms like “Adversarial Attacks in Machine Learning” and "Robustness-Enhancing" emerge in the synthetic sample, similar to the topic of “adversarial detection” from the private sample. It shows that the synthetic sample shifts the topic from privacy to robustness over PE iterations, progressively aligning more closely with the private sample.

### v2

@@ C.12 Aug-PE Convergence under One Private Sample
Private Real deal Mexican food in Chadds Ford, PA. Excellent burritos, tacos, and tamales. I plan to be a regular. Meals served traditionally with sliced radishes, pumpkin seeds and lime. Freshly prepared fillings, onion, cilantro, and red and green sauces. Brightly colored building across from Bavarian Motors. Nothing fancy, but great authentic food.
@@
iter 0 I recently visited this restaurant having heard that they had one of the best taco stands in town. I was not disappointed! The tacos were incredibly delicious and tasted like they were made with fresh ingredients. The portions were generous, and the prices were extremely reasonable. The service was excellent too, with the staff being attentive and accommodating. The ambience of the restaurant was also great, making it a perfect spot for a casual night out. Overall, I had a fantastic experience and would definitely recommend this place to all those who love tacos. A definite 5-star rating!
@@
iter 1 Despite having heard that they had one of the best taco stands in town, I was not disappointed. The tacos were incredibly delicious and tasted like they were made with fresh ingredients. The portions were generous, and the prices were reasonable. The staff was attentive and accommodating. The ambience was great, making it a perfect spot for a casual night out. Overall, I had a fantastic experience and would highly recommend this place to all those craving Mexican cuisine. A solid 5-star rating from me!
@@
iter 2 Despite having heard that they had some of the best Mexican food in town, the crowded parking lot stands in contrast. However, that was not the case. The food tasted authentic and fresh, with the guacamole and salsa being particularly noteworthy. The portions were generous, and the prices were reasonable. The staff was attentive and accommodating. The ambience was great, making it a perfect spot for a date night or a family dinner. Overall, I had a fantastic experience and would highly recommend this
@@
iter 3 Many had heard that they had amazing Mexican food in town, but the crowded parking lot stands in contrast. However, once you step inside, the experience is incredible. The food is authentic, and the guacamole and salsa being particularly noteworthy. Prices were as well quite reasonable. The staff was attentive and accommodating. The ambiance was great, making it perfect for a date night or family dinner. I had a fantastic experience and highly recommend the restaurant to anyone seeking quality Mexican cuisine.
@@
iter 4 Many locals had heard that they had the best Mexican food in town and the hype stands in its truth. However, upon stepping into the experience, it was incredible. The food was fresh, flavorful, and authentic with the guacamole and salsa being particularly noteworthy. The portions were well-sized and satisfying. The staff was attentive and accommodating. The ambiance was cozy and intimate, making it perfect for a romantic night out or casual dinner with friends. I had a great time and highly
@@
iter 5 Many locals had raved about the best authentic Mexican food in town and they were not exaggerating. Walking in to the restaurant was refreshing. The food was fresh, the guacamole and salsa were noteworthy. The margaritas were strong and satisfying. The staff were attentive and accommodating. The ambiance was cozy and intimate, making it perfect for a romantic dinner with loved ones. The prices were great and the portions were generous. The fajitas were sizzling and the tacos were packed with flavor. Overall, this
@@
iter 6 Many locals have raved about the best authentic Mexican food in town and they were not exaggerating. Walking in the restaurant was refreshing. The decor was colorful, lively and festive. The margarita was perfectly balanced and satisfying. The waiter was attentive and accommodating. The ambiance was cozy and intimate, making it perfect for a romantic dinner. The chips and salsa were packed with flavor. Overall, the food was great and the portions were generous. The fajitas were sizzling and the chicken was tender. The guacam
@@
iter 7 Many locals consistently rave about the amazing authentic Mexican cuisine and we were not disappointed. The flavors were refreshing, the decor colorful and lively, and the margaritas satisfying. The waiter was attentive and accommodating. The outdoor seating was perfect for a romantic dinner and the chips and salsa were packed with flavor. Overall, the food was top-notch and the portions were generous. The chicken enchiladas and guacamole were particularly noteworthy. It was a great experience and we highly recommend this restaurant to anyone looking for a delicious meal and a
@@
iter 8 Many foodies rave about the amazing authentic Mexican cuisine and they were not wrong. The flavors were richly robust, colorful and enticing, and the margaritas were top-notch. The enchiladas and guacamole were particularly outstanding. The service was attentive and accommodating and the outdoor seating was perfect for a leisurely dinner and people-watching. From the chips and salsa to the flavorful entrees, the food was superb with generous portions. The menu offers a large selection of delectable options which makes it easy to recommend to anyone looking for a satisfying meal
@@
iter 9 Many foodies have raved about the amazing authentic Mexican cuisine. The flavors were bold and delicious. The tacos were rich and flavorful, with enticing toppings and freshly made guacamole. The service was attentive and accommodating, and the outdoor seating was perfect for a warm summer evening. From the crispy chips and flavorful salsa to the perfectly spiced dishes, the food was generous and satisfying. The menu offers a large selection of vegetarian and meat options which makes it easy to please anyone’s taste buds. Overall, the experience was excellent and highly recommended for people looking for an authentic Mexican dining experience.
@@
Private The authors propose a new statistic, called Expected Perturbation Score (EPS), for adversarial detection. Based on EPS, the authors develop a Maximum Mean Difference (MMD) metric to measure the difference between test samples and natural samples, and further propose an EPS-based adversarial detection method (EPS-AD). Sufficient theoretical analysis and extensive experiments demonstrate the correctness and effectiveness of the proposed method. Strength: 1. This paper is easy to read. 2. The authors propose a new statistical method, called Expected Perturbation Score (EPS), which is able to obtain enough information to identify adversarial examples with only one example after various perturbations. 3. Sufficient theoretical analysis is performed to demonstrate that EPS is able to simulate the difference between the two distributions under mild conditions. Furthermore, extensive experimental results demonstrate the superiority of the proposed EPS-AD. 4. The proposed EPS will be an effective statistic in many applications, such as out-of-distribution detection and anomaly detection. Weaknesses: This paper is easy to read and interesting but still has some minor issues, please refer to weaknesses.
@@
iter 0 Review Example: Overall, this paper provides a comprehensive discussion on privacy issues related to machine learning applications. The authors discuss the concept of differential privacy as a potential solution to protect individual privacy while still maintaining the accuracy of the model. They also provide examples of implementation and evaluation of privacy-preserving techniques using real-world datasets. Strengths: - The paper provides a thorough explanation of differential privacy and how it can be utilized in various machine learning techniques. - The use of real-world datasets to demonstrate implementation and evaluation of privacy techniques adds credibility to the work. - The paper highlights the importance of privacy in machine learning and provides readers with potential solutions to protect individual privacy. Weaknesses: - More evaluation of the privacy techniques could be provided to better understand their effectiveness. - The paper focuses primarily on differential privacy and does not explore other privacy-preserving techniques. - The practical implementation of the techniques may be complex for those without a strong understanding of machine learning and data analysis. In summary, this paper is a strong contribution to the area of social aspects of machine learning, specifically in the field of privacy. It analyzes the concept of differential privacy, its implementation, and evaluation to protect individual privacy in machine learning models. Despite some limitations, the paper contributes significantly to the field and is recommended for acceptance.
@@
iter 1 In their paper ’Privacy in Machine Learning: An Overview’, the authors provide a comprehensive discussion on privacy issues in machine learning. The authors discuss the concept of differential privacy as a potential solution for individual privacy while maintaining the utility of the data. They also provide examples and analysis of privacy-preserving techniques on real-world datasets such as the Stanford and MNIST datasets. The thorough explanation of differential privacy and how it can be utilized in various machine learning techniques adds credibility to the paper. The paper highlights the importance of privacy in machine learning and provides a potential solution to individual privacy. Weaknesses: More evaluation of the privacy techniques could be provided to better understand their effectiveness. The paper primarily focuses on differential privacy and does not explore other privacy-preserving techniques. The technical implementation of the techniques may be difficult for those without a strong understanding of machine learning and statistical analysis. Overall, this paper contributes significantly to the area of social aspects of machine learning, and it is recommended for acceptance as a good paper.
@@
iter 2 In this paper, titled ’Policy Machine Learning: An Overview of the Discussion on Privacy Issues in Machine Learning’, the authors discuss privacy as a potential concern for individual privacy while maintaining the integrity of the data. They also provide examples of privacy-preserving techniques on real-world data from Stanford and MNIST. The thorough exploration of differential privacy can be utilized effectively which adds credibility to the paper. The paper emphasizes the importance of privacy in machine learning and provides a valuable contribution to the field. Weaknesses include the evaluation of techniques to be used to assess their effectiveness. The paper focuses on privacy issues and does not explore fairness-preserving methods. With its contribution to the social aspects of machine learning and statistical analysis, the paper is recommended with a rating of 8 as a good paper.
@@
iter 3 In their research paper, Inference Attack Policy Machine Learning: An Interpretable and Almost True Framework for Predictive Analytics, the authors highlight potential concerns for individual privacy while discussing the importance of privacy in machine learning. They also provide examples of how sensitive data from ImageNet and MNIST datasets can be utilized effectively while ensuring thorough differential privacy which adds credibility to the paper. The research emphasizes the importance of interpretability in machine learning, making a valuable contribution to the field of social aspects of machine learning. We recommend including case studies of how interpretability can be used to assess their effectiveness. The paper also outlines how it does not explore fairness and ethics methods. With this contribution to the field of machine learning and statistical modeling, the authors provide a valuable framework for policy inference attack in machine learning.
@@
iter 4 In their research paper, ’Inference Attacks in Machine Learning: An Interpretability and Almost Interpretability Framework and its Application to Privacy and Analytics’, the authors highlight the need for protecting sensitive data in machine learning. They provide examples of sensitive data from ImageNet and NIST datasets, emphasizing the importance of being thorough in privacy protection to ensure credibility to their research. The paper stresses the importance of interpretability in machine learning. By making a valuable contribution to this field, it provides case studies of how interpretability can be used to assess the effectiveness of machine learning models. The paper outlines various approaches to exploring fairness, transparency, and ethics in machine learning. The results of the study contribute to the need for a comprehensive policy to prevent inference attacks in machine learning.
@@
iter 5 In our research paper, entitled ’Adversarial Attacks in Machine Learning: An Interpretability-Almost-Explainability Framework and its Application to Private Data Analysis’, the authors emphasize the need for protecting sensitive data in machine learning. They provide examples by using data from Inet and MNIST dataset, and address the importance of privacy to ensure the credibility of the results. The paper is well-written and well-structured, making a valuable contribution to the field. Additionally, it highlights the importance of interpretability to enhance the effectiveness of machine learning models. The paper also focuses on fairness, transparency, and ethics in machine learning and the study presents a comprehensive analysis in adversarial attacks. We highly recommend accepting this good paper.
@@
iter 6 In our research paper titled ’Adversarial Attacks in Machine Learning: An Interpretable and Robustness-Enhancing Framework and Empirical Data Analysis’, the authors emphasize the significance of interpretability in machine learning. They provide a comprehensive approach using Integrated Gradients and M-Taylor expansions, to address the challenges and ensure the robustness of results. The paper is well-written, making valuable contributions to the field, and emphasizes the importance of interpretability to enhance the effectiveness of machine learning. Moreover, the study presents a comprehensive approach in defending against adversarial attacks. Therefore, I recommend accepting this good paper.
@@
In this section, we only use one private example in Alg. 1 to generate one synthetic sample. We qualitatively examine if the synthetic sample from Aug-PE increasingly resembles this specific private sample over the PE iterations. This offers a clearer illustration of Aug-PE’s convergence behavior. Specifically, at each iteration, we generate $K$ variations for the current synthetic sample, use the private sample to identify and vote for its nearest synthetic sample based on their embeddings, and select the nearest synthetic sample for the next iteration. Tb. 36 and Tb. 37 show the generations results from GPT-3.5 under one Yelp private sample and one OpenReview private sample, respectively.
@@
As shown Tb. 36, after the voting, the selected synthetic sample relates to the term “taco”, a word present in the private example. By the second iteration, the synthetic sample includes the term “Mexican food”, which aligns with the central theme of the private example. By the fifth iteration, the phrase “authentic Mexican food” surfaces in the synthetic sample, resonating with phrases like “real deal Mexican food” and “great authentic food” from the private example. This demonstrates that the synthetic sample increasingly aligns with the private sample as the iterations progress.
@@
In the OpenReview example presented in Tb. 36, we note that the initial synthetic sample at iteration 0 pertains to the privacy aspects of machine learning, whereas the private sample focuses on adversarial detection and robustness. As the iterations progress, by iteration 4, the topic of synthetic sample shifts to “inference attack in machine learning”, which aligns with the robustness theme of the private sample. By the fifth iteration, terms like “Adversarial Attacks in Machine Learning” and "Robustness-Enhancing" emerge in the synthetic sample, similar to the topic of “adversarial detection” from the private sample. It shows that the synthetic sample shifts the topic from privacy to robustness over PE iterations, progressively aligning more closely with the private sample.